OpenStudy (anonymous):

When 16 litres of water are added to a holding tank that is 1/2 full the tank is then 2/3 full. how many litres can the tank hold ? ( please show ur work)

OpenStudy (anonymous):

let x liters can hold so x/2+16=2x/3 16=x/6 x=96 liters

OpenStudy (anonymous):

thanks ! :3